# a possible strange quark star using the Cherenkov Telescope Array Observatory

### v1

@@ 4 Conclusions
Strangelet formation is more likely during transitions to the CFL phase than the 2SC phase due to the former’s flavor symmetry and enhanced stability. Mixed phases between the 2SC and CFL states may serve as nucleation sites for strangelets, though their stability is compromised by surface and Coulomb effects. Magnetic fields and chiral symmetry restoration significantly influence phase stability and strangelet production. Observational signatures in neutron stars, such as r-mode damping, and particle production in heavy-ion collisions offer promising avenues for detection.
If HESS J1731-347 contains strange quark matter, as in a strange star scenario, the CTA could detect or constrain monochromatic gamma-ray lines (e.g., from $\widetilde{S}\widetilde{S}\to\gamma\gamma$), yielding limits competitive with dark matter searches due to similar annihilation signatures. The CTA’s advanced sensitivity, resolution, and sky coverage make it an ideal instrument for constraining strangelet spectral lines in HESS J1731-347, potentially setting stringent limits on flux and number density. The Galactic Plane Survey and multi-messenger approaches will further enhance these searches, necessitating new theoretical models to predict specific strangelet signatures and advance our understanding of exotic matter in the universe.

### v2

@@ 4 Conclusions
The 2SC phase, characterized by paired up and down quarks, is a superconducting state that may exist in the dense cores of neutron stars. The CFL phase, involving up, down, and strange quarks, exhibits enhanced flavor symmetry and stability, associated with SQM. The transition from the 2SC to the CFL phase can promote strangelet nucleation, potentially transforming a neutron star into a strange quark star, with stability influenced by magnetic fields, chiral symmetry restoration, and surface/Coulomb effects.
The CTA is optimized to detect VHE gamma rays, ranging from tens of GeV to hundreds of TeV. In the context of the compact object HESS J1731-347, which may harbor SQM in a strange star scenario, the CTA has the potential to detect or constrain gamma-ray lines resulting from strangelet annihilation. Additionally, the CTA’s sensitivity to such annihilation signatures enables it to set competitive limits in dark matter searches, as these signatures are analogous to those expected from dark matter particle annihilation.
The CTA’s advanced sensitivity, resolution, and sky coverage make it an ideal instrument for constraining strangelet spectral lines in HESS J1731-347, potentially setting stringent limits on flux and number density. The Galactic Plane Survey and multi-messenger approaches will further enhance these searches, necessitating new theoretical models to predict specific strangelet signatures and advance our understanding of exotic matter in the universe.